# Monte Carlo simulations of relativistic shock breakout from a stellar wind

### v1

@@ Abstract
[…]
@@ 1 Introduction
Shock breakout governs the initial electromagnetic emission from various cosmic explosions, including supernovae and gamma-ray bursts. During breakout, photons that were initially trapped within the transition region of a radiation-mediated shock (RMS) are released (see, e.g., Waxman & Katz, 2017; Levinson & Nakar, 2020). A detailed understanding of RMS structure is therefore essential for interpreting the properties of breakout emission.
The shock velocity at the breakout can range from sub-relativistic to mildly relativistic, and in extreme cases even ultra- relativistic, as anticipated, e.g., in energetic explosions of a compact progenitor, or highly collimated explosions such as low-luminosity gamma-ray bursts (LLGRBs). When the velocity of RMS is slow ($\beta_{u}\lesssim 0.05$), where $\beta_{u}$ denotes the shock velocity normalized by the speed of light $c$, thermal equilibrium is achieved throughout the shock. Additionally, the diffusion approximation can be applied to solve radiation transfer within the shock. These aspects significantly simplify the problem, and the physics of slow RMSs has been extensively studied and is well established in the literature, particularly in the regime well before breakout where the shock resides in a highly optically thick region ($\tau\gg\beta_{u}^{-1}$), and thus the steady-state assumption holds (e.g., Zel’dovich & Raizer, 1967; Weaver, 1976; Blandford & Payne, 1981).
@@
The studies outlined above focus on the steady-state profiles of shocks in which photons are trapped. While these studies offer valuable insights, they cannot be directly applied to the breakout phase, during which photons escape from the shock. In particular, shock breakout occurring in a region with a sharp density drop, such as at the edge of a stellar envelope or the outermost layers of an extended envelope, is inherently an abrupt phenomenon. Accurately capturing the resulting emission requires time-dependent modeling. In the case of slow and fast shocks, substantial progress has been made through both analytical approaches (e.g., Chevalier, 1992; Nakar & Sari, 2010; Katz et al., 2012; Irwin & Hotokezaka, 2024) and radiation hydrodynamics simulations, which adopt the diffusion approximation for radiative transfer with varying simplifications across studies (e.g., Klein & Chevalier, 1978; Ensman & Burrows, 1992; Blinnikov et al., 2000; Tolstov et al., 2013; Sapir et al., 2013; Sapir & Halbertal, 2014). These radiation-hydrodynamics-based simulations have explored the evolution of the shock structure during breakout and the resulting emission. In contrast, studies of relativistic shock breakout remain limited and are largely restricted to analytical treatments (e.g., Nakar & Sari, 2012; Faran & Sari, 2023). These studies suggest that the breakout flash is triggered when the shocked ejecta expands sufficiently and its temperature decreases to approximately $50~{}{\rm keV}$. At this point, previously generated pairs annihilate, allowing photons that had been trapped by pair opacity to escape and produce the observable breakout emission. However, a fully self-consistent calculation of this process has not yet been conducted.  [^1] The only ab initio simulation of relativistic RMS breakout to date is by Lundman & Beloborodov (2021), which focuses on the photon-rich regime but neglects pair production, limiting its applicability.
@@
Shock breakout can also occur in shallowly decaying density regions, such as the optically thick stellar wind surrounding a progenitor star (Ofek et al., 2014; Gal-Yam et al., 2014), representing an important class of breakout events alongside those occurring at sharp density drop regions. Numerous studies have investigated shock breakout in stellar winds, particularly in the sub-relativistic regime (e.g., Balberg & Loeb, 2011; Chevalier & Irwin, 2011; Svirski et al., 2012; Svirski & Nakar, 2014; Khatami & Kasen, 2024). A key difference, also highlighted in the present study, is that in shallowly decaying density regions, the breakout proceeds relatively gradually, provided that the shock-crossing time is comparable to (not significantly shorter than) the timescale over which the upstream conditions evolve during breakout. This is in contrast to breakouts at the stellar edge, where the upstream density drops steeply and the breakout is more abrupt. Taking advantage of this more gradual transition, photon-starved shock breakout in winds has been modeled both analytically and numerically in recent years.  [^2] However, a recent high-resolution radiation hydrodynamics study by Wasserman et al. (2025) suggests that steady-state treatments may be invalid under these conditions. Ioka et al. (2019) developed a simple analytic model for RMS breakout in the fast Newtonian regime ($\beta_{u}\lesssim 0.3$), incorporating radiative losses under the diffusion approximation. This was followed by numerical simulations by Ito et al. (2020b) (hereafter ILN20b), which also assumed steady-state conditions but employed full radiative transfer. These simulations showed good agreement with the analytic model for shock velocities of $\beta_{u}=0.1$ and $\beta_{u}\lesssim 0.25$. At higher velocities, approaching the mildly relativistic regime ($\beta_{u}\approx 0.5$), the simulations revealed vigorous pair production and the formation of a collisionless subshock, both of which substantially alter the shock structure and emergent spectrum.
@@
Similarly, under the steady-state assumption, wind breakout in the highly relativistic regime has been analytically investigated by Granot et al. (2018) (hereafter GNL18). Their analysis indicate that vigorous pair production within the shock transition layer generates sufficient self-generated opacity to sustain the shock to be radiation-mediated during the breakout phase. Unlike relativistic breakouts at sharp edges, where the breakout flash originates abruptly from a shocked layer with optical depth near unity, the breakout emission arises from the gradual leakage of photons from the shock front, which advances over orders of magnitude in radius as the fraction of photons escaping from the shock increases. The model introduced in GNL18, however, has some important limitations. First, it is based on assumptions that are applicable only in highly relativistic regimes; hence, the shock profile cannot be properly extended to the downstream region. Second, as a result, the model was expected to fail in describing mildly relativistic shocks. Third, the model cannot be used to compute the spectrum. Given these limitations, a more comprehensive study is required to deepen our understanding of relativistic RMS, particularly in the presence of photon escape. To achieve this, we employ a numerical approach developed in Ito et al. (2020b), which allows us to extend the analysis beyond the highly relativistic regime and investigate shocks with mildly relativistic velocities. Specifically, we examine cases with $\Gamma_{u}=2$, $6$, and $10$. Our numerical approach enables us to construct a complete shock structure, spanning from the upstream (escaping) boundary to the far downstream region, and to evaluate the spectral properties of the emitted radiation. Contrary to naive expectations, our results show that mildly relativistic shocks with $\Gamma_{u}=2$ can be well reproduced by the analytical model, regardless of the amount of photon escape from the shock. In contrast, for higher Lorentz factors ($\Gamma_{u}=6$ and $10$), while a shock without photon escape shows excellent agreement with the analytical model, the model breaks down as photon escape from the shock becomes significant. This behavior is discussed in detail in this paper.
@@
This paper is organized as follows. In Section 2, we outline the numerical methods and the simulation setup. Section 3 presents the calculated structure of RMS, while Section 4 discusses the spectral characteristics of photons escaping from RMS. The comparison between our numerical results and the analytical model of GNL18 is presented in Section 5, and its implications for the observable properties of relativistic shock breakout from a stellar wind are discussed in Section 6. Finally, we conclude in Section 7. Throughout the paper, the subscript $u$ and $d$ refer to the physical quantities at the far upstream and far downstream regions of the shock, respectively.
@@ 2 Monte-Carlo simulations of leaking RRMS
The numerical method used in the current study is outlined in ILN20b, and we provide a brief summary of the methodology here. The code is an updated version of the one used in ILN20a, which focused on the steady-state structure of RMS with an optically thick upstream that prevents photons from escaping from the shock (referred to hereafter as ’infinite shock’). The updated code incorporates radiative losses through the upstream boundary, enabling simulations of RRMS with finite optical depths upstream, as anticipated during the breakout phase. Similar modifications have been applied previously to simulations of fast Newtonian ($\beta_{u}=0.1$ and $0.25$) and mildly relativistic ($\beta_{u}=0.5$) shock breakouts, as reported in ILN20b. In this study, we extend our analysis to the relativistic regime, specifically considering cases with $\Gamma_{u}=2$, $6$, and $10$.
[…]
@@ 3 THE STRUCTURE OF RRMS WITH ESCAPE
Figures 1 - 3 display the shock structures derived from simulations for upstream Lorentz factors $\Gamma_{u}=2$, $6$, and $10$ across a wide range of escape fractions $f_{esc}$. As depicted in the figure, the shock width significantly reduces with increasing $f_{esc}$. This result contrasts with the sub-relativistic shocks (Ioka et al., 2019, ILN20b), where the shock width is mostly consistent with the diffusion length $\tau_{*}\sim\beta_{u}^{-1}$, regardless of the value of $f_{esc}$. This is due to the fact that relativistic shocks have an ability to adjust its opacity by generating pairs (GNL18). The current simulation finds that a minute amount of escape $f_{esc}\lesssim 0.01$ is sufficient to appreciably narrow the shock width, particularly when the Lorentz factor of the shock is high. This feature is consistent with the analytical model of GNL18, which suggests that finite shocks deviate markedly from the infinite shock configuration when the escape fraction exceeds $\sim 1/\Gamma_{u}^{2}$. However, it should be noted that our results also reveal quantitative differences in the shock structure from that of the analytical model at high escape fractions, as will be discussed in Section 5.
[…]
@@
As mentioned earlier, the subshock, characterized by a discontinuous jump in the velocity and temperature, is an inherent feature of relativistic shocks regardless of the escape fraction. [^6] As reported in ILN20b, a subshock is not present in the sub-relativistic regime ($\beta_{u}=0.1$ and $0.25$). In the mildly relativistic regime, where $\beta_{u}=0.5$, the subshock is absent in the case of an infinite shock; however, it manifests in finite shocks for finite shocks ($f_{esc}>0$). In our simulation, the jumps between pre- and - post-subshock velocity and temperature are connected by the Rankine-Hugoniot conditions, under the assumption that the bulk plasma is isolated from the radiation (Ito et al., 2018). The strength of the subshock increases with $f_{esc}$, as observed in ILN20b for mildly relativistic shocks with $\beta_{u}=0.5$.  [^7] To be precise, in our iterative calculations, the subshock strength does not converge perfectly but fluctuates within a range of roughly $10$-$20\%$ in the jump of 4-velocity $\delta(\Gamma\beta)$. Consequently, the subshock might occasionally appear slightly stronger for smaller escape fractions. Specifically, the sequence of subshock strength depicted in Figure 1 ($\Gamma_{u}=2$) is not directly proportional to the escape fraction in the range of $f_{esc}=0\%$ to $6\%$. The fluctuations observed in the current simulations might suggest that a perfectly steady profile is unattainable, indicating that the subshock feature could be inherently dynamical. However, exploring this aspect further is beyond the scope of the current study. For instance, in shocks with $\Gamma_{u}=10$, the 4-velocity at the immediate upstream of the subshock, $\Gamma_{u,sub}\beta_{u,sub}$, is approximately $0.84$, $0.85$, $1.08$, $1.4$, $2.25$, and $4.35$ for $f_{esc}\simeq 0$, $0.41$, $2.5$, $5.9$, $14$, and $41~{}\%$, respectively. This results in a fraction of the total kinetic energy being dissipated at the subshock, which we evaluate as the reduction in kinetic energy flux of the plasma across the subshock, normalized by the incoming kinetic energy flux at the upstream boundary expressed as $\delta F_{kin,sub}/F_{b}=[1+(n_{\pm}/n)_{sub}m_{e}/m_{p}](\Gamma_{u,sub}-\Gamma_{d,sub})/(\Gamma_{u}-1)$, to increase as $4.8$, $4.8$, $8.8$, $14$, $25$, and $48\%$ for the corresponding values of $f_{esc}$. [^8] For $\Gamma_{u}=2$ ($6$), the reduction of kinetic energy across the subshock $\delta F_{kin,sub}/F_{b}$ is roughly $15\%$ ($45\%$) for the simulation with the highest escape fraction of $f_{esc}=35\%$ ($46\%$). Here, $\Gamma_{d,sub}$ denotes the Lorentz factor immediately downstream of the subshock, and $(n_{\pm}/n)_{sub}$ indicates the pair-to-baryon number density ratio at the subshock. The result implies that the role of radiation mediation in the shock weakens as $f_{esc}$ increases, and a significant amount of energy is dissipated by the plasma interactions at the subshock in finite shocks with high escape fractions. This is in contrast with the sub-relativistic shocks, where subshocks do not appear even at a high escape fraction ($f_{esc}\sim 70\%$ for $\beta\leq 0.25$; ILN20b).
@@
As mentioned above, the subshock also leads to a jump in the temperature, due to the dissipation of plasma kinetic energy, manifesting as a discontinuous spike in Figures 1 - 3. Due to the proportionality of the subshock strength with the escape fraction, simulations with larger escape fractions exhibit a more pronounced temperature spike. While the temperature in most regions is predominantly determined by the Compton temperature, where the net heating and cooling of the pairs balance, this discontinuous jump causes the temperature to deviate temporarily from equilibrium. However, the plasma quickly cools via inverse Compton scattering and rejoins the Compton equilibrium temperature within a pair-loaded optical depth of a few. The equilibrium temperature subsequent to this spike falls within the range of $kT\sim 100$-$200$ keV, irrespective of the values of $f_{esc}$ or $\Gamma_{u}$. This regulation of temperature is attributed to the vigorous pair production, which functions as a thermostat, preventing the temperature from substantially exceeding the pair production threshold, a characteristic property of RRMS well-discussed in the literature (see Levinson & Nakar, 2020, and references therein). As a result, except for the initial rise at the upstream boundary and the spike at the subshock, the temperature exhibits a similar profile among different escape fractions.

### v2

@@ Abstract
We present Monte Carlo simulations of relativistic radiation-mediated shocks (RRMS) in the photon-starved regime, incorporating photon escape from the upstream region—characterized by the escape fraction, $f_{\rm esc}$—under a steady-state assumption. These simulations, performed for shock Lorentz factors $\Gamma_{u}=2$, $3.5$, 66, $10$, and $15$, are applicable to RRMS breakouts in shallowly declining density profiles such as stellar winds. We find that vigorous pair production acts as a thermostat, regulating the downstream temperature to $\sim 100$-$200~{\rm keV}$, largely independent of $f_{\rm esc}$. A subshock forms and strengthens with increasing $f_{\rm esc}$. The escaping spectra peak at $E_{p}\approx 300$-$600~{\rm keV}$ in the shock frame and deviate from a Wien distribution, exhibiting low-energy flattening ($f_{\nu}\propto\nu^{0}$) due to free-free emission and high-energy extensions caused by inverse Compton scattering from subshock-heated pairs. While an earlier analytical model reproduces the velocity structure well at $\Gamma_{u}=2$, it significantly overestimates the shock width at higher Lorentz factors, particularly for $f_{\rm esc}\gtrsim$ a few $\%$. Based on this finding, we provide updated predictions for breakout observables in wind environments for $\Gamma_{u}\gtrsim 6$. Notably, the duration of the relativistic breakout becomes largely insensitive to the explosion energy and ejecta mass, typically exceeding analytical predictions by orders of magnitude and capable of producing a $\sim$300 s flash of MeV photons with a radiated energy of $\sim 10^{50}$erg for an energetic explosion yielding $\Gamma_{bo}\sim 6$. We also discuss limitations of our modelling assumptions and their implications for the predicted breakout observables.
keywords:
shock waves – radiative transfer – radiation mechanisms: thermal – radiation mechanisms: non-thermal – methods: numerical – stars: winds, outflows
@@ 1 Introduction
Shock breakout governs the initial electromagnetic emission from various cosmic explosions, including supernovae and gamma-ray bursts. During breakout, photons that were initially trapped within the transition region of a radiation-mediated shock (RMS) are released (see, e.g., Waxman and Katz, 2017; Levinson and Nakar, 2020). A detailed understanding of RMS structure is therefore essential for interpreting the properties of breakout emission.
The shock velocity at the breakout can range from sub-relativistic to mildly relativistic, and in extreme cases even ultra- relativistic, as anticipated, e.g., in energetic explosions of a compact progenitor, or highly collimated explosions such as low-luminosity gamma-ray bursts (LLGRBs). When the velocity of RMS is slow ($\beta_{u}\lesssim 0.05$), where $\beta_{u}$ denotes the shock velocity normalized by the speed of light $c$, thermal equilibrium is achieved throughout the shock. Additionally, the diffusion approximation can be applied to solve radiation transfer within the shock. These aspects significantly simplify the problem, and the physics of slow RMSs has been extensively studied and is well established in the literature, particularly in the regime well before breakout where the shock resides in a highly optically thick region ($\tau\gg\beta_{u}^{-1}$), and thus the steady-state assumption holds (e.g., Zel’dovich and Raizer, 1967; Weaver, 1976; Blandford and Payne, 1981).
@@
The studies outlined above focus on the steady-state profiles of shocks in which photons are trapped. While these studies offer valuable insights, they cannot be directly applied to the breakout phase, during which photons escape from the shock. In particular, shock breakout occurring in a region with a sharp density drop, such as at the edge of a stellar envelope or the outermost layers of an extended envelope, is inherently an abrupt phenomenon. Accurately capturing the resulting emission requires time-dependent modeling. In the case of slow and fast shocks, substantial progress has been made through both analytical approaches (e.g., Chevalier, 1992; Nakar and Sari, 2010; Katz et al., 2012; Irwin and Hotokezaka, 2024) and radiation hydrodynamics simulations, which adopt the diffusion approximation for radiative transfer with varying simplifications across studies (e.g., Klein and Chevalier, 1978; Ensman and Burrows, 1992; Blinnikov et al., 2000; Tolstov et al., 2013; Sapir et al., 2013; Sapir and Halbertal, 2014). These radiation-hydrodynamics-based simulations have explored the evolution of the shock structure during breakout and the resulting emission. In contrast, studies of relativistic shock breakout remain limited and are largely restricted to analytical treatments (e.g., Nakar and Sari, 2012; Faran and Sari, 2023). These studies suggest that the breakout flash is triggered when the shocked ejecta expands sufficiently and its temperature decreases to approximately $50~{\rm keV}$. At this point, previously generated pairs annihilate, allowing photons that had been trapped by pair opacity to escape and produce the observable breakout emission. However, a fully self-consistent calculation of this process has not yet been conducted.  [^1] The only ab initio simulation of relativistic RMS breakout to date is by Lundman and Beloborodov (2021), which focuses on the photon-rich regime but neglects pair production, limiting its applicability.
@@
Shock breakout can also occur in shallowly decaying density regions, such as the optically thick stellar wind surrounding a progenitor star (Ofek et al., 2014; Gal-Yam et al., 2014), representing an important class of breakout events alongside those occurring at sharp density drop regions. Numerous studies have investigated shock breakout in stellar winds, particularly in the sub-relativistic regime (e.g., Balberg and Loeb, 2011; Chevalier and Irwin, 2011; Svirski et al., 2012; Svirski and Nakar, 2014; Khatami and Kasen, 2024). A key difference, also highlighted in the present study, is that in shallowly decaying density regions, the breakout proceeds relatively gradually, provided that the shock-crossing time is comparable to (not significantly shorter than) the timescale over which the upstream conditions evolve during breakout. This is in contrast to breakouts at the stellar edge, where the upstream density drops steeply and the breakout is more abrupt. Taking advantage of this more gradual transition, photon-starved shock breakout in winds has been modeled both analytically and numerically in recent years.  [^2] However, a recent high-resolution radiation hydrodynamics study by Wasserman et al. (2025) suggests that steady-state treatments may be invalid under these conditions. Ioka et al. (2019) developed a simple analytic model for RMS breakout in the fast Newtonian regime ($\beta_{u}\lesssim 0.3$), incorporating radiative losses under the diffusion approximation. This was followed by numerical simulations by Ito et al. (2020b) (hereafter ILN20b), which also assumed steady-state conditions but employed full radiative transfer. These simulations showed good agreement with the analytic model for shock velocities of $\beta_{u}=0.1$ and $\beta_{u}\lesssim 0.25$. At higher velocities, approaching the mildly relativistic regime ($\beta_{u}\approx 0.5$), the simulations revealed vigorous pair production and the formation of a collisionless subshock, both of which substantially alter the shock structure and emergent spectrum.
@@
Similarly, under the steady-state assumption, wind breakout in the highly relativistic regime has been analytically investigated by Granot et al. (2018) (hereafter GNL18). Their analysis indicates that vigorous pair production within the shock transition layer generates sufficient self-generated opacity to sustain the shock to be radiation-mediated during the breakout phase. Unlike relativistic breakouts at sharp edges, where the breakout flash originates abruptly from a shocked layer with optical depth near unity, the breakout emission arises from the gradual leakage of photons from the shock front, which advances over orders of magnitude in radius as the fraction of photons escaping from the shock increases. The model introduced in GNL18, however, has some important limitations. First, it is based on assumptions that are applicable only in highly relativistic regimes; hence, the shock profile cannot be properly extended to the downstream region. Second, as a result, the model was expected to fail in describing mildly relativistic shocks. Third, the model cannot be used to compute the spectrum. Given these limitations, a more comprehensive study is required to deepen our understanding of relativistic RMS, particularly in the presence of photon escape. To achieve this, we employ a numerical approach developed in Ito et al. (2020b), which allows us to extend the analysis beyond the highly relativistic regime and investigate shocks with mildly relativistic velocities. Specifically, we examine cases with $\Gamma_{u}=2$, $3.5$ 66, $10$, and $15$. Our numerical approach enables us to construct a complete shock structure, spanning from the upstream (escaping) boundary to the far downstream region, and to evaluate the spectral properties of the emitted radiation. Contrary to naive expectations, our results show that mildly relativistic shocks with $\Gamma_{u}=2$ can be well reproduced by the analytical model, regardless of the amount of photon escape from the shock. In contrast, for higher Lorentz factors ($\Gamma_{u}\geq 3.5$),while a shock without photon escape shows excellent agreement with the analytical model, the model breaks down as photon escape from the shock becomes significant. This behavior is discussed in detail in this paper. We note that both the GNL18 model and our numerical study adopt a steady-state, planar treatment of the shock structure; a definitive validation of this approximation would require fully time-dependent simulations, which are beyond the scope of the present work.
@@
This paper is organized as follows. In Section 2, we outline the numerical methods and the simulation setup. Section 3 presents the calculated structure of RMS, while Section 4 discusses the spectral characteristics of photons escaping from RMS. The comparison between our numerical results and the analytical model of GNL18 is presented in Section 5, and its implications for the observable properties of relativistic shock breakout from a stellar wind are discussed in Section 6. We discuss the limitations of our modelling assumptions, including the steady-state planar treatment, in Section 7. Finally, we conclude in Section 8. Throughout the paper, the subscript $u$ and $d$ refer to the physical quantities at the far upstream and far downstream regions of the shock, respectively.
@@ 2 Monte-Carlo simulations of leaking RRMS
The numerical method used in the current study is outlined in ILN20b, and we provide a brief summary of the methodology here. The code is an updated version of the one used in ILN20a, which focused on the steady-state structure of RMS with an optically thick upstream that prevents photons from escaping from the shock (referred to hereafter as ’infinite shock’). The updated code incorporates radiative losses through the upstream boundary, enabling simulations of RRMS with finite optical depths upstream, as anticipated during the breakout phase. Similar modifications have been applied previously to simulations of fast Newtonian ($\beta_{u}=0.1$ and $0.25$) and mildly relativistic ($\beta_{u}=0.5$) shock breakouts, as reported in ILN20b. In this study, we extend our analysis to the relativistic regime, specifically considering cases with $\Gamma_{u}=2$, $3.5$, 66, $10$, and $15$.
In the simulations, the pair-loaded optical depth upstream of the subshock, which is defined as $\tau_{u*}=\int_{0}^{z_{u}}\Gamma(n+n_{\pm})\sigma_{T}dz$, serves as an input parameter that dictates the amount of photons escaping the upstream boundary. Here, $\sigma_{T}$ denotes the Thomson cross-section. The term $n$ refers to the density of the electrons advected from upstream, while $n_{\pm}$ represents the density of electron-positron pairs produced via photon-photon collisions. In this study, we focus on RMS formed in a proton-electron plasma; therefore, the density of advected electrons is equal to that of the baryons (protons). Throughout the paper, we use $n$ to represent the density of both the advected electrons and protons. The range of integration extends from the subshock at $z=0$ to the upstream boundary at $z_{u}$ (<0<0). [^3] Since the spatial coordinate $z$ is aligned with the flow direction, $\tau_{*u}$ is defined as a negative value. Note that, as found in ILN20a for infinite shocks, the formation of a subshock is an inherent feature in relativistic shocks, also in cases of finite shocks.
@@ 3 THE STRUCTURE OF RRMS WITH ESCAPE
Figures 1 - 4 display the shock structures derived from simulations for upstream Lorentz factors $\Gamma_{u}=2$, $3.5$, 66, and $10$ across five values of the escape fraction $f_{esc}$ spanning 0 to $\lesssim 0.5$. We omit a description of the $\Gamma_{u}=15$ shock structure here and the corresponding spectra in Section 4 because we performed only two runs at $f_{esc}=0.14$ and $0.4$, which do not cover the $f_{\rm esc}$ dependence from 0 to high values. We note, however, that these runs show no qualitative differences from the high-$f_{esc}$ cases at $\Gamma_{u}=10$. As described in Section 5, we use these two runs to examine how the shock width depends on $\Gamma_{u}$ at high $f_{\rm esc}$. As depicted in the figure, the shock width significantly reduces with increasing $f_{esc}$. This result contrasts with the sub-relativistic shocks (Ioka et al., 2019, ILN20b), where the shock width is mostly consistent with the diffusion length $\tau_{*}\sim\beta_{u}^{-1}$, regardless of the value of $f_{esc}$. This is due to the fact that relativistic shocks have an ability to adjust its opacity by generating pairs (GNL18). The current simulation finds that a minute amount of escape $f_{esc}\lesssim 0.01$ is sufficient to appreciably narrow the shock width, particularly when the Lorentz factor of the shock is high. This feature is consistent with the analytical model of GNL18, which suggests that finite shocks deviate markedly from the infinite shock configuration when the escape fraction exceeds $\sim 1/\Gamma_{u}^{2}$. However, it should be noted that our results also reveal quantitative differences in the shock structure from that of the analytical model at high escape fractions, as will be discussed in Section 5.
Figure 1: Profiles of 4-velocity (top), temperature (middle), and pair-to-baryon ratio (bottom) as functions of the pair-loaded Thomson optical depth $\tau_{*}=\int\Gamma(n+n_{\pm})\sigma_{T}dz$ for the simulations with an upstream Lorentz factor of $\Gamma_{u}=2$. The profiles are represented by solid lines in green, orange, purple, blue, and red, each indicating the results from finite shocks with different escape fractions as detailed in the legend. For comparative purposes, the structure of an infinite shock is depicted with a black line. It is noted that for infinite shocks, the upstream boundary extends beyond the left edge of the figure, whereas the leftmost points of the finite shock profiles mark their respective upstream boundaries. A discontinuous jump in both the 4-velocity and temperature profiles at $\tau_{*}=0$ reflects the presence of a subshock.
Figure 2: Same as Figure 1, but for $\Gamma_{u}=3.5$.
@@
As mentioned earlier, the subshock, characterized by a discontinuous jump in the velocity and temperature, is an inherent feature of relativistic shocks regardless of the escape fraction. [^6] As reported in ILN20b, a subshock is not present in the sub-relativistic regime ($\beta_{u}=0.1$ and $0.25$). In the mildly relativistic regime, where $\beta_{u}=0.5$, the subshock is absent in the case of an infinite shock; however, it manifests in finite shocks ($f_{esc}>0$). In our simulation, the jumps between pre- and post-subshock velocity and temperature are connected by the Rankine-Hugoniot conditions, under the assumption that the bulk plasma is isolated from the radiation (Ito et al., 2018). The strength of the subshock increases with $f_{esc}$, as observed in ILN20b for mildly relativistic shocks with $\beta_{u}=0.5$.  [^7] To be precise, in our iterative calculations, the subshock strength does not converge perfectly but fluctuates within a range of roughly $10$-$20\%$ in the jump of 4-velocity $\delta(\Gamma\beta)$. Consequently, the subshock might occasionally appear slightly stronger for smaller escape fractions. Specifically, the sequence of subshock strength depicted in Figure 1 ($\Gamma_{u}=2$) is not directly proportional to the escape fraction in the range of $f_{esc}=0\%$ to $6\%$. The fluctuations observed in the current simulations might suggest that a perfectly steady profile is unattainable, indicating that the subshock feature could be inherently dynamical. However, exploring this aspect further is beyond the scope of the current study. For instance, in shocks with $\Gamma_{u}=10$, the 4-velocity at the immediate upstream of the subshock, $\Gamma_{u,sub}\beta_{u,sub}$, is approximately $0.84$, $0.85$, $1.08$, $1.4$, $2.25$, and $4.35$ for $f_{esc}\simeq 0$, $0.41$, $2.5$, $5.9$, $14$, and $41~\%$, respectively. This results in a fraction of the total kinetic energy being dissipated at the subshock, which we evaluate as the reduction in kinetic energy flux of the plasma across the subshock, normalized by the incoming kinetic energy flux at the upstream boundary expressed as $\delta F_{kin,sub}/F_{b}=[1+(n_{\pm}/n)_{sub}m_{e}/m_{p}](\Gamma_{u,sub}-\Gamma_{d,sub})/(\Gamma_{u}-1)$, to increase as $4.8$, $4.8$, $8.8$, $14$, $25$, and $48\%$ for the corresponding values of $f_{esc}$. [^8] For $\Gamma_{u}=2$ (66), the reduction of kinetic energy across the subshock $\delta F_{kin,sub}/F_{b}$ is roughly $15\%$ ($45\%$) for the simulation with the highest escape fraction of $f_{esc}=35\%$ ($46\%$). Here, $\Gamma_{d,sub}$ denotes the Lorentz factor immediately downstream of the subshock, and $(n_{\pm}/n)_{sub}$ indicates the pair-to-baryon number density ratio at the subshock. The result implies that the role of radiation mediation in the shock weakens as $f_{esc}$ increases, and a significant amount of energy is dissipated by the plasma interactions at the subshock in finite shocks with high escape fractions. This is in contrast with the sub-relativistic shocks, where subshocks do not appear even at a high escape fraction ($f_{esc}\sim 70\%$ for $\beta\leq 0.25$; ILN20b).
@@
As mentioned above, the subshock also leads to a jump in the temperature, due to the dissipation of plasma kinetic energy, manifesting as a discontinuous spike in Figures 1 - 4. Due to the proportionality of the subshock strength with the escape fraction, simulations with larger escape fractions exhibit a more pronounced temperature spike. While the temperature in most regions is predominantly determined by the Compton temperature, where the net heating and cooling of the pairs balance, this discontinuous jump causes the temperature to deviate temporarily from equilibrium. However, the plasma quickly cools via inverse Compton scattering and rejoins the Compton equilibrium temperature within a pair-loaded optical depth of a few. The equilibrium temperature subsequent to this spike falls within the range of $kT\sim 100$-$200$ keV, irrespective of the values of $f_{esc}$ or $\Gamma_{u}$. This regulation of temperature is attributed to the vigorous pair production, which functions as a thermostat, preventing the temperature from substantially exceeding the pair production threshold, a characteristic property of RRMS well-discussed in the literature (see Levinson and Nakar, 2020, and references therein). As a result, except for the initial rise at the upstream boundary and the spike at the subshock, the temperature exhibits a similar profile among different escape fractions.